OpenStudy (anonymous):

hookies law. the distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. if the distance is 50 cm when the weight is 7kg what is the distance when the weight is 5kg.

OpenStudy (anonymous):

F = -kd

OpenStudy (anonymous):

or W = -kd

OpenStudy (anonymous):

or W = kd

OpenStudy (whynot):

F=kx mg=kx--1 m1=7kg x1=0.5m m1g=kx1 k=m1g/x1 k=137.2 m2g=kx2 5kg*9.8m/s^2=137.2kg/s^2*x2 x2=0.357142 m x2=35.7

OpenStudy (whynot):

35.7cm

OpenStudy (jamesj):

And for the record, it's not Hookies Law or alas even Hookers Law, but Hooke's Law.